First of all, here comes the examples from the float package, the rotated versions of these will follow on the next pages:

Program 0.1 The first program. This hasn’t got anything to do with the style but is included as an example. Note the ruled float style.

#include <stdio.h>
int main(int argc, char **argv)
{
  int i;
  for (i = 0; i < argc; ++i)
          printf("argv[%d] = %s\n", i, argv[i]);
  return 0;
}

\floatstyle{ruled}
\newfloat{Program}{tbp}{lop}[section]

… loads o’ stuff …

\begin{Program}
\begin{verbatim}

… program text …

\end{verbatim}

\caption{… caption … }

\end{Program}

Example 0.1: This is another silly floating Example.

\begin{matrix}n & \binom{n}{0} & \binom{n}{1} & \binom{n}{2} & \binom{n}{3} & \binom{n}{4} & \binom{n}{5} & \binom{n}{6} & \binom{n}{7} \\ 0 & 1 \\ 1 & 1 & 1 \\ 2 & 1 & 2 & 1 \\ 3 & 1 & 3 & 3 & 1 \\ 4 & 1 & 4 & 6 & 4 & 1 \\ 5 & 1 & 5 & 10 & 10 & 5 & 1 \\ 6 & 1 & 6 & 15 & 20 & 15 & 6 & 1 \\ 7 & 1 & 7 & 21 & 35 & 35 & 21 & 7 & 1 \end{matrix}

Table 1: Pascal’s triangle. This is a re-styled L^{A}TEX table.

Program 0.2 The first program. This hasn’t got anything to do with the style but is included as an example. Note the ruled float style.

#include <stdio.h>
int main(int argc, char **argv)
{
       int i;
       for (i = 0; i < argc; ++i)
               printf("argv[%d] = %s\n", i, argv[i]);
       return 0;
}

2

\floatstyle{ruled}
\newfloat{Program}{tbp}{lop}[section]

… loads o’ stuff …

\begin{Program}
\begin{verbatim}

… program text …

\end{verbatim}

\caption{… caption … }

\end{Program}

Example 0.2: This is another silly floating Example.

3

\begin{matrix}n & \binom{n}{0} & \binom{n}{1} & \binom{n}{2} & \binom{n}{3} & \binom{n}{4} & \binom{n}{5} & \binom{n}{6} & \binom{n}{7} \\ 0 & 1 \\ 1 & 1 & 1 \\ 2 & 1 & 2 & 1 \\ 3 & 1 & 3 & 3 & 1 \\ 4 & 1 & 4 & 6 & 4 & 1 \\ 5 & 1 & 5 & 10 & 10 & 5 & 1 \\ 6 & 1 & 6 & 15 & 20 & 15 & 6 & 1 \\ 7 & 1 & 7 & 21 & 35 & 35 & 21 & 7 & 1 \end{matrix}

Table 2: Pascal’s triangle. This is a re-styled L^{A}TEX table.